OpenStudy (anonymous):

(4w2 – 7w – 6) – (8w2 + 2w – 3) A. –4w2 – 9w – 3 B. 12w2 + 9w + 3 C. –4w2 – 5w – 9 D. 12w2 – 5w – 9

OpenStudy (aripotta):

4x2 - 8w2 = -4w2 -7w - 2w = -9w -6 - (-3) = -3

OpenStudy (aripotta):

do you understand? it would be a

OpenStudy (anonymous):

@AriPotta, No i dont understand

OpenStudy (aripotta):